OpenStudy (timaashorty):

How far apart are the foci of an ellipse with a major axis of 34 feet and a minor axis of 16 feet? a. 15 ft b. 12 ft c. 9 ft d. 30 ft

OpenStudy (mertsj):

If the major axis is 34, what is the value of "a" ?

OpenStudy (jdoe0001):

\(\bf c=\sqrt{a^2-b^2} \\ \quad \\ c=\textit{distance from the center to either foci}\)

OpenStudy (jdoe0001):

so as @Mertsj said, get the value for the "a" component and the "b" component or minor axis, and solve for "c"

OpenStudy (jdoe0001):



OpenStudy (mertsj):

Are you there, Tim?

OpenStudy (timaashorty):

Yes I'm here now, sorry my dad called me. Give me a few secs to go over what you guys wrote c;

OpenStudy (timaashorty):

\[c=\sqrt{34^2-16^2}\] Do I first solve this?

OpenStudy (timaashorty):

\[c=\sqrt{1156 - 256}\] \[c= 30\]

OpenStudy (timaashorty):

Is that it ?

OpenStudy (timaashorty):

@jdoe0001 or @Mertsj ?

OpenStudy (mertsj):

That is why I asked you to identify the value of a. Look at the picture. the length of the major axis is 2a.

OpenStudy (mertsj):

a=17 b=8

OpenStudy (timaashorty):

how did you get that?

OpenStudy (timaashorty):

\[c=\sqrt{289-64}\] \[c= \sqrt 205\] \[c= 14.32\]

OpenStudy (mertsj):

Now you can answer the question the problem asks.

OpenStudy (timaashorty):

I'm still not positive 15 ft?

OpenStudy (mertsj):

T

OpenStudy (mertsj):

The focal points lie on the major axis. Each focal point is c units from the center. So the distance between the foci is 2c

OpenStudy (timaashorty):

2* 14.32 = 28.64 which is approximetly 30 ft?

OpenStudy (timaashorty):

is that correct?

OpenStudy (mertsj):

yes

OpenStudy (timaashorty):

Thank you!

OpenStudy (mertsj):

I think you make a mistake in your calculation.

OpenStudy (mertsj):

289-64=225

OpenStudy (mertsj):

so c = 15 and 2c = 30